OpenStudy (anonymous):

lim x->0 ((cos t) -1)/ sin t

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ \cos x -1 }{ \sin t }\]

OpenStudy (freckles):

try multiply top's conjugate on top and bottom

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{\cos(x)-1}{\sin(x)} \cdot \frac{\cos(x)+1}{\cos(x)+1} \\ =\lim_{x \rightarrow 0}\frac{\cos^2(x)-1}{\sin(x)(\cos(x)+1)}\] and what is 1-cos^2(x)=? (think trig identity)

OpenStudy (anonymous):

just change to neg sin^2(x)

OpenStudy (freckles):

ok great you should see something cancel from top and bottom now

OpenStudy (anonymous):

so neg sinx/(cos x +1)

OpenStudy (anonymous):

so 0

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{-\sin^2(x)}{\sin(x)(\cos(x)+1)}=\lim_{x \rightarrow 0}\frac{-\sin(x)}{\cos(x)+1}\] yep :)

OpenStudy (anonymous):

tyty

OpenStudy (freckles):

np :)